OpenStudy (anonymous):

Solve the linear inequality 8-7/6x >9 -

OpenStudy (zzr0ck3r):

8-(7/6)x > 9?

OpenStudy (anonymous):

multiply both sides of the inequality by 6

OpenStudy (zzr0ck3r):

-(7/6)x>1 x< -6/7

OpenStudy (anonymous):

Okay! Thanks so much guys. I'll let you all know if I get confused.

OpenStudy (anonymous):

Confused!

OpenStudy (anonymous):

I am multiply 6 by 8-7 and 9?

OpenStudy (anonymous):

the 6 will cancel out on the left side

OpenStudy (anonymous):

Yes, I got that far. Next I multiplied 6 by 9 and got 54

OpenStudy (anonymous):

ya thats correct

OpenStudy (anonymous):

Is that all? There aren't anymore steps to this equation?

OpenStudy (anonymous):

8-7

OpenStudy (zzr0ck3r):

did i write the equation corectly up at the top?

OpenStudy (anonymous):

think you did but my mind is trying to rush through the equation cause they are all due in 30 minutes.

OpenStudy (zzr0ck3r):

\[8-\frac{ 7 }{ 6 }x>9\]

OpenStudy (zzr0ck3r):

?

OpenStudy (anonymous):

> -

OpenStudy (anonymous):

otherwise you are correct

OpenStudy (zzr0ck3r):

subtract 8 from both sides and you are left with \[-\frac{ 7 }{ 6 }x>1\] then multiply both sides by the recipicle of -7/6 \[-\frac{ 6 }{ 7 }*-\frac{ 7 }{ 6 }x \le -\frac{ 6 }{ 7 }\]

OpenStudy (zzr0ck3r):

we flip the sign when deviding with a negative or multiplyingso we are now left with \[x \le -\frac{ 6 }{ 7 }\]

OpenStudy (anonymous):

Okay. thank you!